The sum of the ages of Kwasi and Ama is 40 years. Kwasi is 10 years older than Ama. How old are they

Answer by rothauserc(4718)

let k be Kwasi's age and a be Ama's age
:
1) k + a = 40
:
2) k = a + 10
:
substitute for k in equation 1
:
a + 10 + a = 40
:
2a = 30
:
a = 15, then k = 15 + 10 = 25
:
*********************************************
Kwasi is 25 years old and Ama is 15 years old
*********************************************
: